# Role of flow topology in wind-driven wildfire propagation

### v1

@@ 2.3 Numerical methods
In this section, we present our approach to solving the wildfire model using the finite difference method (FDM), with the additional advantage of leveraging CUDA support to facilitate simulations on GPUs. We implement an upwind compact scheme for the advective term to maintain its upwinding properties and employ central difference schemes for the diffusion term to accurately mirror its isotropic nature. Given the inherent numerical stiffness of the wildfire equation, we utilize the Implicit-Explicit Runge-Kutta (IMEX-RK) method for temporal integration, carefully distinguishing between stiff and non-stiff terms within the newly non-dimensionalized wildfire equation. Additionally, we incorporate a localized artificial diffusivity model to mitigate numerical instabilities, which is discussed below. The validation of our developed FDM solver is presented in Appendix B.
@@ 2.3.1 Upwind compact FDM scheme
The compact schemes with spectral-like resolution were initially introduced by Lele (1992) within the FDM framework. While this scheme exhibits superior performance for periodic boundary conditions due to its implicit nature, it tends to manifest numerical instabilities such as anti-diffusion and anti-dissipation near the domain boundaries when applied to non-periodic problems. Furthermore, optimization of the compact scheme for fluid flow problems was conducted by Sengupta et al. (2003) through global spectral analysis, where the near-boundary numerical instabilities for non-periodic problems were mitigated by the development of higher-order explicit upwind boundary closures, effectively resolving length scales without the need for numerical filtering or damping. The optimized upwind compact scheme (OUCS2) utilized for computing the first derivative advection term at interior grid points is expressed as follows
@@ 2.3.2 IMEX-RK method
The fuel reaction term in the wildfire equation imposes a more severe time-step restriction than the diffusion term, which makes the long-time simulation laborious when one uses a fully explicit time integration method, and it becomes intractable for a fully implicit time integration scheme where one has to solve a system of nonlinear equations. On the other hand, the splitting technique may be helpful for solving the wildfire transport equation with the dynamics at different time scales, provided spatio-temporal balancing is ensured over time. The IMEX-RK schemes provide a viable strategy to integrate stiff and non-stiff terms of an equation, simultaneously, using an implicit and explicit temporal scheme, respectively. In this work, we employed the robust L-stable IMEX-RK3 scheme outlined by Pareschi and Russo (2005) for temporal integration across all wildfire simulations investigated. Notably, Bürger et al. (2020) introduced the IMEX-RK temporal scheme for the wildfire equation while identifying stiff and non-stiff terms for the first time, and we extended their algorithm to accommodate the new non-dimensional form of the equation, presented in Appendix A. We highlight the stiff ($\overline{T}$) and non-stiff non-dimensional temperature ($\widehat{T}$) components within our newly derived non-dimensional wildfire equation, given as follows:
@@ 3 Scaling analysis
First, we perform a scaling analysis to understand the influence of the dimensionless parameters on combustion dynamics within the newly derived non-dimensional wildfire equation. Below, we introduce characteristic scales that, under the specified assumptions, transform the spatial derivative terms in the differential equations into an algebraic form. This analytical process provides valuable insights into the relative significance of various terms and parameters, thereby unveiling the critical determinants governing transient wildfire behavior. The chosen characteristic scales for scaling analysis are as follows
@@ 4.1 Steady wind velocity: A saddle-type fixed point
Figure 2: The two wind patterns considered in this study. (a) Steady wind topology showcasing both stable (blue) and unstable (red) manifolds, with a saddle point positioned at the origin, superimposed with velocity streamlines. (b) Unsteady double gyre wind velocity at $\tau=0$, where the dividing streamline (blue) shows the boundary between the two vortices.
We investigate the transient behavior of wildfires within steady wind velocity fields. By utilizing a saddle-type fixed point in the wind velocity topology, characterized by both attraction and repulsion, we analyze the complex dynamics of fire spread, examining how wildfires respond to varying wind patterns across the spatial domain and providing valuable insights into their behavior under diverse conditions. The selected saddle-type wind topology, with stable and unstable eigenvalues of $-1$ and $1$, is constructed by formulating wind velocity components $u=\xi$ and $v=-\eta$, ensuring the conservation of mass. As depicted in Fig. 2a, the saddle point is located at the origin $(0,0)$. The stable (repelling) and unstable (attracting) manifolds are also shown. Consistent parameters are used across all simulations to ensure coherence and comparability, including a square domain length of $4$, spanning coordinates from $-2$ to $2$ with $\beta=1$ everywhere across a $256\times 256$ grid. A fixed time step of $\Delta\tau=10^{-7}$ is used for time integration, and additional parameters include $q=1$, $\alpha=10^{-3}$, and $\overline{T}_{\text{pc}}=3$. The initial Heaviside square firefront temperature is uniformly set to $\overline{T}=31$, positioned at coordinates $(-1.05,1.05)$ with a square side length of $0.1$, and a normalized maximum wind speed of $|\overrightarrow{\mathbf{w}}|=1$. To maintain consistency across simulations, a stabilization parameter of $C\mu=0.80$ is used, although this LAD model may not be essential for scenarios with moderate $\Phi$ and $Da$ values. To characterize the transient behavior of the initialized firefront, we track the instantaneous locations of four firefronts over time: the top firefront ($F_{T}^{Y}$) advecting along the positive $\eta$ direction, the bottom firefront ($F_{B}^{Y}$) along the negative $\eta$ direction, the right firefront ($F_{R}^{X}$) along the positive $\xi$ direction, and the left firefront ($F_{L}^{X}$) along the negative $\xi$ direction. The corresponding group velocities are indicated by $V_{T}^{Y}$, $V_{B}^{Y}$, $V_{R}^{X}$, and $V_{L}^{Y}$, respectively. This tracking method provides a detailed understanding of the firefronts’ dynamics and progression under varying convection-diffusion-reaction processes.
@@ 4.1.1 Role of Diffusion–Da number
Figure 3: Spatio-temporal evolution of the Heaviside firefront under a saddle-type steady wind velocity field over a uniformly distributed fuel bed, computed across a range of Da values $\in(1,10,10^{2},10^{3})$, with $\Phi=1.0$ and $\epsilon=0.03$. (a) Fuel and temperature fields at various time instants are displayed in the top and bottom rows, respectively, for $Da=10^{3}$, with streamlines superimposed at $\tau=0.01$. At $\tau=0.15$, the firefront locations are marked by white strips, labeled as the top firefront ($F_{T}^{Y}$), bottom firefront ($F_{B}^{Y}$), right firefront ($F_{R}^{X}$), and left firefront ($F_{L}^{X}$), with superscripts indicating their advection direction. (b) The instantaneous spatial advection of the firefronts and their corresponding time-averaged group velocities are shown, depicted in the top and bottom rows, respectively. The dotted horizontal line represents the asymptotically converged group velocity of the firefronts at an infinite $Da$ number.
@@
First, we focus on understanding the role of diffusion on the transient behavior of wildfires by varying the $Da$ number across a range of values $(1,10,10^{2},10^{3})$, while keeping $\Phi=1.0$ and $\epsilon=0.03$ constant. The simulation results for $Da=10^{3}$ are presented in Fig. 3a, providing insights into the spatio-temporal evolution of the firefront over the uniform fuel bed. At $\tau=0.15$, the positions of the four tracked firefronts are visually represented by white strips overlaying the fuel bed, progressively advancing along their respective directions over time, while the provided fuel bed undergoes complete gradual consumption, contributing to the fire’s progression without being influenced by the manifolds. In the top row of Fig. 3b, the instantaneous positions of these firefronts are depicted for the chosen distinct Da values, indicating the group velocity through their slopes. The group velocity is calculated instantaneously for the given Da and firefronts, followed by temporal averaging. The corresponding temporally averaged group velocities of the firefronts are illustrated in the bottom row of Fig. 3b for different Da values.
@@
Interestingly, the consistent linear progression in the advection of all four firefronts is noted, even as the firefront advances beyond the saddle point in the given wind field. Furthermore, the observed decrease in slope values with increasing Da values across all cases suggests an exponential reduction in group velocity with each increment of Da, eventually converging towards an asymptotic value. As a result, an exponential fit regression, expressed as $ae^{bDa}+c$, is employed to quantify the relationship between the firefront’s group velocity and the Da number. Here, the coefficients $a$, $b$, and $c$ represent the amplitude, the exponential decay of group velocity, and the asymptotically converged group velocity at infinite Da number, respectively. The regression results presented in Table 1 show that the four firefronts converge to distinct asymptotic group velocities, reflecting the influence of the applied wind topology. The results emphasize the significant role of the diffusion process in shaping the transient behavior of wildfires, primarily by modulating the magnitude of their firefront’s group velocity while leaving the direction of wildfire propagation unaltered.
@@
     $c$      $3.673$      $3.704$      $3.318$      $3.524$
Table 1: Characterizing the time-averaged group velocities of the firefronts for the initialized Heaviside firefront under a saddle-type steady wind velocity field over a uniformly distributed fuel bed, using exponential fitting, for different $Da\in(1,10,10^{2},10^{3})$, $\Phi=1.0$, and $\epsilon=0.03$.
@@ 4.1.2 Role of Convection–$\Phi$ number
Figure 4: Spatio-temporal evolution of the Heaviside firefront under a saddle-type steady wind velocity field over a uniformly distributed fuel bed, computed across a range of $\Phi$ values $\in(0.1,0.05,0.01)$, with $Da=10^{3}$ and $\epsilon=0.03$. (a) Fuel and temperature fields at various time instants, with the top and bottom rows, respectively, computed for $\Phi=0.05$. Streamlines are superimposed at $\tau=0.1$ and $\tau=0.6$. At $\tau=0.60$, the dotted white curve highlights the local neutral curve, indicating the stalled top and bottom firefronts over the fuel bed. (b) Instantaneous spatial advection of the firefronts and their corresponding time-averaged group velocities, depicted in the top and bottom rows, respectively. The red dots highlight the group velocity of stalled firefronts along their respective advecting directions.
To investigate the influence of convection on wildfire propagation, $\Phi$ is varied across values $(0.1,0.05,0.01)$, indicative of increasing wind velocity with decreasing $\Phi$. Da number of $10^{3}$ is selected to minimize the diffusion dominance along with $\epsilon=0.03$. The spatio-temporal evolution of the Heaviside firefront at various time instants is depicted in Fig. 4a, illustrating both the fuel and temperature fields in the top and bottom rows, respectively, computed for $\Phi=0.05$. A significant departure from the last section arises here, where the observed fuel bed does not undergo complete consumption. Instead, the advection of the firefronts and the corresponding fuel consumption rate are influenced by imposed stable/unstable manifolds. The attracting streamlines propel both the left and right firefronts forward, effectively encouraging complete fuel consumption along their trajectory. Conversely, the repelling streamlines exert a decelerating effect on the top and bottom firefronts, inducing a phenomenon akin to “stalling”, where the fire’s progression is halted, leaving unburnt fuel remnants. This intriguing observation prompts the identification of the spatial location of the stalled top and bottom firefronts, marked by a white dotted curve at $\tau=0.6$ over the fuel bed in Fig. 4a. Furthermore, at this spatial location, one observes that the convection, diffusion, and reaction contributions balances with each other (the right-hand side of Eq. 16), hence both the temperature temporal rate $(\partial\overline{T}/\partial\tau)$ and fuel burn rate $(\partial\beta/\partial\tau)$ reach zero. This notable equilibrium delineates the spatial boundary of the stalled firefronts as a “local neutral curve”, beyond which the fire’s progression ceases and cooling gradually occurs over time as all the fuel has been consumed.
@@ 4.2 Transient wind velocity: Double gyre flow
We delve into investigating transient wind velocity fields, considering the double gyre flow, a typical benchmark in environmental flows. Double gyre flow is characterized by a pair of two counter-rotating vortices, as shown in Fig. 2b. The double gyre flow is constructed by formulating wind velocity components as
@@
where $f(\xi,\tau)=\lambda\sin(2\pi\Omega\tau)\xi^{2}+[1-2\lambda\sin(2\pi\Omega\tau)]\xi$, with $\lambda$ the perturbation amplitude representing how far the dividing streamline perturbs from the initial position, $\Omega$ the oscillation frequency indicating the rate of this perturbation, and $A_{m}$ the velocity magnitude. The wind oscillation frequency, normalized by the total fuel consumption timescale $\overline{t}_{r}$ (the time required to completely burn all the given fuel), is represented by the Strouhal number $\textit{St}=\overline{t}_{r}\Omega$, quantifying the interaction between fuel reaction and flow oscillation timescales. Consistent parameters across simulations include a rectangular domain of $2\times 1$ along the $\xi$ and $\eta$ axes, a $256\times 128$ grid, $\beta=1$, a time step of $\Delta\tau=10^{-8}$, $q=1$, $\alpha=10^{-3}$, $\overline{T}_{\text{pc}}=3$, and an initial Heaviside square firefront temperature $\overline{T}=31$ at $(1.0,0.5)$ with a side length of $0.1$ along the dividing streamline. Simulations employ a normalized maximum wind speed $|\overrightarrow{\mathbf{w}}|=1$ and LAD coefficient $C\mu=0.75$ to examine transient wildfire dynamics under convection-dominant and low-diffusion scenarios, defined by $\Phi=10^{-3}$, $Da=10^{3}$, and $\epsilon=0.03$. Given the transient wind topology, instantaneous FTLE fields were computed both forward and backward in time over an integration period of $20$, allowing for the extraction of repelling and attracting Lagrangian coherent structures (rLCS and aLCS) at a given time $\tau$.

### v2

@@ 2.3 Numerical methods
In this section, we present our approach to solving the physics-based CDR wildfire model using the finite difference method (FDM), with the additional advantage of leveraging CUDA support to facilitate simulations on GPUs. We implement an upwind compact scheme for the advective term to maintain its upwinding properties and employ central difference schemes for the diffusion term to accurately mirror its isotropic nature. Given the inherent numerical stiffness of the wildfire equation, we utilize the Implicit-Explicit Runge-Kutta (IMEX-RK) method for temporal integration, carefully distinguishing between stiff and non-stiff terms within the newly non-dimensionalized wildfire equation. Additionally, we incorporate a localized artificial diffusivity model to mitigate numerical instabilities, which is discussed below. The detailed FDM solver algorithm and the corresponding validation of the implementation are provided in Appendix A and Appendix B, respectively. Additionally, Appendix C presents a comparative analysis of three wildfire modeling strategies—a semi-empirical approach, the newly derived CDR model, and a fully coupled CFD framework—focusing on their rate of spread prediction capabilities.
@@ 2.3.1 Upwind compact FDM scheme
The compact schemes with spectral-like resolution were initially introduced by Lele (1992) within the FDM framework. While this scheme exhibits superior performance for periodic boundary conditions due to its implicit nature, it tends to manifest numerical instabilities such as anti-diffusion and anti-dissipation near the domain boundaries when applied to non-periodic problems. Furthermore, optimization of the compact scheme for fluid flow problems was conducted by Sengupta et al. (2003) through global spectral analysis, where the near-boundary numerical instabilities for non-periodic problems were mitigated by the development of higher-order explicit upwind boundary closures, effectively resolving length scales without the need for numerical filtering or damping. As proposed in Sengupta et al. (2003), the optimized fifth-order upwind compact scheme (OUCS2) is utilized for computing the first derivative advection term at interior grid points is expressed as follows
@@ 2.3.2 IMEX-RK method
The fuel reaction term in the wildfire equation imposes a more severe time-step restriction than the diffusion term, which makes the long-time simulation laborious when one uses a fully explicit time integration method, and it becomes intractable for a fully implicit time integration scheme where one has to solve a system of nonlinear equations. On the other hand, the splitting technique may be helpful for solving the wildfire transport equation with the dynamics at different time scales, provided spatio-temporal balancing is ensured over time. The IMEX-RK schemes provide a viable strategy to integrate stiff and non-stiff terms of an equation, simultaneously, using an implicit and explicit temporal scheme, respectively. In this work, we employed the robust L-stable IMEX-RK3 scheme outlined by Pareschi and Russo (2005) for temporal integration across all wildfire simulations investigated. Notably, Bürger et al. (2020) introduced the IMEX-RK temporal scheme for the wildfire equation while identifying stiff and non-stiff terms for the first time, and we extended their algorithm to accommodate the new non-dimensional form of the equation. We highlight the stiff ($\overline{T}$) and non-stiff non-dimensional temperature ($\widehat{T}$) components within our newly derived non-dimensional wildfire equation, given as follows:
@@ 3 Scaling analysis
First, we perform a scaling analysis to understand the influence of the dimensionless parameters on the underlying transient reactive flow dynamics within the newly derived non-dimensional physics-based CDR wildfire equation. Below, we introduce characteristic scales that, under the specified assumptions, transform the spatial derivative terms in the differential equations into an algebraic form. This analytical process provides valuable insights into the relative significance of various terms and parameters, thereby unveiling the critical determinants governing transient wildfire behavior within the physics-based CDR model. The chosen characteristic scales for scaling analysis are as follows
@@ 4.1 Steady wind velocity: A saddle-type fixed point
Figure 2: The two wind patterns considered in this study. (a) Steady wind topology showcasing both stable (blue) and unstable (red) manifolds, with a saddle point positioned at the origin, superimposed with velocity streamlines. (b) Unsteady double gyre wind velocity at $\tau=0$, where the dividing streamline (blue) shows the boundary between the two vortices. The vertical dividing streamline undergoes translation motion along the $\xi$ axis, oscillating about the value of $1$, with a given amplitude and frequency over time.
We investigate the reactive flow dynamics of the simplified physics-based CDR wildfire model within the context of steady wind velocity fields. By utilizing a saddle-type fixed point in the wind velocity topology, characterized by both attraction and repulsion, we analyze the complex dynamics of fire spread, examining how wildfires respond to varying wind patterns across the spatial domain and providing valuable insights into their behavior under diverse conditions. The selected saddle-type wind topology, with stable and unstable eigenvalues of $-1$ and $1$, is constructed by formulating wind velocity components $u=\xi$ and $v=-\eta$, ensuring the conservation of mass. As depicted in Fig. 2a, the saddle point is located at the origin $(0,0)$. The stable (repelling) and unstable (attracting) manifolds are also shown. Consistent parameters are used across all simulations to ensure coherence and comparability, including a square domain length of $4$, spanning coordinates from $-2$ to $2$ with $\beta=1$ everywhere across a $256\times 256$ grid. A fixed time step of $\Delta\tau=10^{-7}$ is used for time integration, and additional parameters include $q=1$, $\alpha=10^{-3}$, $\overline{\kappa}=0.1$, and $\overline{T}_{\text{pc}}=3$. The initial Heaviside square firefront temperature is uniformly set to $\overline{T}=31$, positioned at coordinates $(-1.05,1.05)$ with a square side length of $0.1$, and a normalized maximum wind speed of $|\overrightarrow{\mathbf{w}}|=1$. The corresponding dimensional initial fire temperatures are $T_{\text{pc}}=327$ K ($\overline{T}_{\text{pc}}=3$) and $T=527$ K ($\overline{T}=31$), for $\epsilon=0.03$ and ambient temperature $T_{\infty}=300$ K, which are sufficient to initiate the flammable exothermic process, as noted by Serón et al. (2005); Sudhakar and Kolar (2011); Vogiatzoglou et al. (2024). To maintain consistency across simulations, a stabilization parameter of $C\mu=0.80$ is used, although this LAD model may not be essential for scenarios with moderate $\Phi$ and $Da$ values. To characterize the transient behavior of the initialized firefront, we track the instantaneous locations of four firefronts over time: the top firefront ($F_{T}^{Y}$) advecting along the positive $\eta$ direction, the bottom firefront ($F_{B}^{Y}$) along the negative $\eta$ direction, the right firefront ($F_{R}^{X}$) along the positive $\xi$ direction, and the left firefront ($F_{L}^{X}$) along the negative $\xi$ direction. The corresponding group velocities are indicated by $V_{T}^{Y}$, $V_{B}^{Y}$, $V_{R}^{X}$, and $V_{L}^{Y}$, respectively. This tracking method provides a detailed understanding of the firefronts’ dynamics and progression under varying convection-diffusion-reaction processes.
@@ 4.1.1 Role of Diffusion–Da number
Figure 3: Spatio-temporal evolution of the Heaviside firefront under a saddle-type steady wind velocity field over a uniformly distributed fuel bed, computed across a range of Da values $\in(1,10,10^{2},10^{3},10^{4},10^{5})$, with $\Phi=1.0$ and $\epsilon=0.03$. (a) Fuel and temperature fields at various time instants are displayed in the top and bottom rows, respectively, for $Da=10^{3}$, with streamlines superimposed at $\tau=0.01$. At $\tau=0.15$, the firefront locations are marked by white strips, labeled as the top firefront ($F_{T}^{Y}$), bottom firefront ($F_{B}^{Y}$), right firefront ($F_{R}^{X}$), and left firefront ($F_{L}^{X}$), with superscripts indicating their movement direction. (b) The instantaneous spatial advection of the firefronts and their corresponding time-averaged group velocities are shown, depicted in the top and bottom rows, respectively. The dotted horizontal line represents the asymptotically converged group velocity of the firefronts at an infinite $Da$ number.
@@
First, we focus on understanding the role of diffusion on the transient behavior of wildfires by varying the $Da$ number across a range of values $(1,10,10^{2},10^{3},10^{4},10^{5})$, while fixing $\Phi=1.0$ and $\epsilon=0.03$ constant. The simulation results for $Da=10^{3}$ are presented in Fig. 3a, providing insights into the spatio-temporal evolution of the firefront over the uniform fuel bed. At $\tau=0.15$, the positions of the four tracked firefronts are visually represented by white strips overlaying the fuel bed, progressively advancing along their respective directions over time, while the provided fuel bed undergoes complete gradual consumption, contributing to the fire’s progression without being influenced by the manifolds. In the top row of Fig. 3b, the instantaneous positions of these firefronts are depicted for the chosen distinct Da values, indicating the group velocity through their slopes. The group velocity is calculated instantaneously for the given Da and firefronts, followed by temporal averaging. The corresponding temporally averaged group velocities of the firefronts are illustrated in the bottom row of Fig. 3b for different Da values.
@@
Interestingly, the consistent linear progression in the advection of all four firefronts is noted, even as the firefront advances beyond the saddle point in the given wind field. Furthermore, the observed decrease in slope values with increasing Da values across all cases suggests an exponential reduction in group velocity with each increment of Da, eventually converging towards an asymptotic value. Additional simulations were performed by extending Da to $10^{9}$; however, the results are not presented here, as asymptotic convergence was observed, particularly when $\textit{Da}>10^{3}$. This indicates that further increasing Da beyond $10^{3}$ (minimizing diffusion further) would not yield a significant influence on the deceleration of the firefronts’ group velocity. Consequently, an exponential fit regression, expressed as $ae^{bDa}+c$, is employed to quantify the relationship between the firefront’s group velocity and the Da number. Here, the coefficients $a$, $b$, and $c$ represent the amplitude, the exponential decay of group velocity, and the asymptotically converged group velocity at infinite Da number, respectively. The regression results presented in Table 1 show that the four firefronts converge to distinct asymptotic group velocities, reflecting the influence of the applied wind topology. The results emphasize the significant role of the diffusion process in shaping the transient behavior of wildfires, primarily by influencing only the “magnitude” of their firefront’s group velocity while leaving the “direction” of wildfire propagation unaltered.
@@
     $c$      $3.619$      $3.704$      $3.318$      $3.524$
Table 1: Characterizing the time-averaged group velocities of the firefronts for the initialized Heaviside firefront under a saddle-type steady wind velocity field over a uniformly distributed fuel bed, using exponential fitting, for different $Da\in(1,10,10^{2},10^{3},10^{4},10^{5})$, $\Phi=1.0$, and $\epsilon=0.03$.
@@ 4.1.2 Role of Convection–$\Phi$ number
Figure 4: Spatio-temporal evolution of the Heaviside firefront under a saddle-type steady wind velocity field over a uniformly distributed fuel bed, computed across a range of $\Phi$ values $\in(100,10,1,0.1,0.05,0.01)$, with $Da=10^{3}$ and $\epsilon=0.03$. (a) Fuel and temperature fields at various time instants, with the top and bottom rows, respectively, computed for $\Phi=0.05$. Streamlines are superimposed at $\tau=0.1$ and $\tau=0.6$. At $\tau=0.60$, the dotted white curve highlights the local neutral curve, indicating the stalled top and bottom firefronts over the fuel bed. (b) Instantaneous spatial advection of the firefronts and their corresponding time-averaged group velocities, depicted in the top and bottom rows, respectively. The red dots highlight the group velocity of stalled firefronts along their respective advecting directions.
To investigate the influence of convection on wildfire propagation, $\Phi$ is varied across values of 100 to 0.01, indicative of increasing wind velocity with decreasing $\Phi$. Da number of $10^{3}$ is selected to minimize the diffusion dominance along with $\epsilon=0.03$. The spatio-temporal evolution of the Heaviside firefront at various time instants is depicted in Fig. 4a, illustrating both the fuel and temperature fields in the top and bottom rows, respectively, computed for $\Phi=0.05$. A significant departure from the last section arises here, where the observed fuel bed does not undergo complete consumption. Instead, the advection of the firefronts and the corresponding fuel consumption rate are influenced by imposed stable/unstable manifolds. The attracting streamlines propel both the left and right firefronts forward, effectively encouraging complete fuel consumption along their trajectory. Conversely, the repelling streamlines exert a decelerating effect on the top and bottom firefronts, inducing a phenomenon akin to “stalling”, where the fire’s progression is halted, leaving unburnt fuel remnants. This intriguing observation prompts the identification of the spatial location of the stalled top and bottom firefronts, marked by a white dotted curve at $\tau=0.6$ over the fuel bed in Fig. 4a. Furthermore, at this spatial location, one observes that the convection, diffusion, and reaction contributions balances with each other (the right-hand side of Eq. 16), hence both the temperature temporal rate $(\partial\overline{T}/\partial\tau)$ and fuel burn rate $(\partial\beta/\partial\tau)$ reach zero. This notable equilibrium delineates the spatial boundary of the stalled firefronts as a “local neutral curve”, beyond which the fire’s progression ceases and cooling gradually occurs over time as all the fuel has been consumed.
@@ 4.2 Transient wind velocity: Double gyre flow
Now we look into the transient reactive flow dynamics of the physics-based CDR model by investigating their complex interaction with unsteady wind velocity topology, using the double gyre flow, a typical benchmark in environmental flows. Double gyre flow is characterized by a pair of two counter-rotating vortices, as shown in Fig. 2b. The double gyre flow is constructed by formulating wind velocity components as
@@
where $f(\xi,\tau)=\lambda\sin(2\pi\Omega\tau)\xi^{2}+[1-2\lambda\sin(2\pi\Omega\tau)]\xi$, with $\lambda$ the perturbation amplitude representing how far the dividing streamline perturbs from the initial position, $\Omega$ the oscillation frequency indicating the rate of this perturbation, and $A_{m}$ the velocity magnitude. The wind oscillation frequency, normalized by the total fuel consumption timescale $\overline{t}_{r}$ (the time required to completely burn all the given fuel), is represented by the Strouhal number $\textit{St}=\overline{t}_{r}\Omega$, quantifying the interaction between fuel reaction and flow oscillation timescales. Consistent parameters across simulations include a rectangular domain of $2\times 1$ along the $\xi$ and $\eta$ axes, a $256\times 128$ grid, $\beta=1$, a time step of $\Delta\tau=10^{-8}$, $q=1$, $\alpha=10^{-3}$, $\overline{T}_{\text{pc}}=3$, $\overline{\kappa}=0.1$, and an initial Heaviside square firefront temperature $\overline{T}=31$ at $(1.0,0.5)$ with a side length of $0.1$ along the dividing streamline. Simulations employ a normalized maximum wind speed $|\overrightarrow{\mathbf{w}}|=1$ and LAD coefficient $C\mu=0.75$ to examine transient wildfire dynamics under convection-dominant and low-diffusion scenarios, defined by $\Phi=10^{-3}$, $Da=10^{3}$, and $\epsilon=0.03$. Given the transient wind topology, instantaneous FTLE fields were computed both forward and backward in time over an integration period of $20$, allowing for the extraction of repelling and attracting Lagrangian coherent structures (rLCS and aLCS) at a given time $\tau$.